OpenStudy (anonymous):

A public library has fiction and nonfiction books in the ratio 3:2. If the library has 360 fiction books ho many nonfiction books does it have?

OpenStudy (king):

240

OpenStudy (anonymous):

240

OpenStudy (ash2326):

fiction/nonfiction=3:2 360/nonfiction=3/2 360*2/3=nonfiction nonfiction=240

OpenStudy (anonymous):

Fict = 3 Nonfict = 2 thus, 360/3 = 120 ( 1 unit is 120) 120*2 = 240